Calculate ΔGorxn given the following information. 4HNO3(g) + 5N2H4(l) → 7N2(g) + 12H2O(l) HNO3(g) N2H4(l) N2(g) H2O(l) ΔGof (kJ/mol) -73.5 149.3 0 -237.1
Answer #1

Answer:. -3297.7 KJ/mol

Calculation and explanation:

Answer Theron is 4 HNO₃ (8) + 5 N₂ H₃ (1) 7N₂ (g) +12 H₂O(1) 460 (K I mail) = 4 *(-735) 6*149-3) (40) 1286-237-) 45. urn = 16